# Spin Theory Based on the Extended Least Action Principle and Information Metrics: Quantization, Entanglement, and Bell Test With Time Delay

### v2

@@ III.3 Subsequent Measurement with a Rotated Stern-Gerlach Apparatus
In Appendix C, we verify that (36) is exactly the same result predicted by standard quantum mechanics.
@@ IV Derivation of the Schrödinger-Pauli Equation
The framework to derive the law of dynamics for an electron in a external magnetic field $B_{z}$, for a cumulative period from $t_{a}$ to $t_{b}$, is similar to that described in Appendix A, except that we need to add a new term in the Lagrangian due to the the interaction of spin and the external magnetic field. Based on the results shown in (16) and (29), the probability density can be written as
@@
$\begin{split}\bar{U}&=-\int p(\theta)\vec{\mu}\cdot\vec{B}d\theta\\ &=\frac{e\hbar}{2m}\int p(\theta)B_{z}\cos(\theta)d\theta,\end{split}$ (38)
@@
$\begin{split}\bar{U}&=\frac{e\hbar}{2m}\int\rho(\mathbf{x},t,\theta)B_{z}\cos(\theta)d\theta d^{3}\mathbf{x}\\ &=\frac{e\hbar}{2m}\sum_{i}\int\sigma_{i}(\mathbf{x},t,\theta_{i})\delta(\theta-\theta_{i})B_{z}\cos(\theta)d\theta d^{3}\mathbf{x}.\end{split}$ (40)
@@
$\begin{split}\bar{U}&=\frac{e\hbar}{2m}\{\int\sigma_{0}(\mathbf{x},t,0)\delta(\theta)\mu B_{z}\cos(\theta)d\theta d^{3}\mathbf{x}\\ &+\int\sigma_{1}(\mathbf{x},t,\pi)\delta(\theta-\pi)\mu B_{z}\cos(\theta)d\theta d^{3}\mathbf{x}\}\\ &=\frac{e\hbar}{2m}\{\int\sigma_{0}(\mathbf{x},t,0)\mu B_{z}d^{3}\mathbf{x}-\int\sigma_{1}(\mathbf{x},t,\pi)\mu B_{z}d^{3}\mathbf{x}\}.\end{split}$ (41)
Here we run into a problem. In the Stern-Gerlach experiment, it is confirmed that electrons with spin-up follow a different trajectory from that of electrons with spin-down. An electron with either spin-up or spin-down will follow separated laws of dynamics. We cannot simply add the two terms in (41) together. Instead, a more proper notation should treat the probability density as a two-component vector. Thus,
$\bar{U}=\frac{e\hbar}{2m}\int B_{z}\left(\begin{array}[]{c}\sigma_{0}(d^{3}\mathbf{x},t,0)\\ -\sigma_{1}(d^{3}\mathbf{x},t,\pi)\end{array}\right)dx:=\left(\begin{array}[]{c}\bar{U}_{+}\\ \bar{U}_{-}\end{array}\right).$ (42)
For further simplification of notation, denote $\rho_{+}(x,t)=\sigma_{0}(x,t,0)$ and $\rho_{-}(x,t)=\sigma_{0}(x,t,0)$. They should satisfy the normalization condition
$\int[\rho_{+}(d^{3}\mathbf{x},t)+\rho_{-}(d^{3}\mathbf{x},t)]dx=1.$ (43)
First we will derive the law of dynamics for the electron with spin up. Without considering the contribution from the spin magnetic interaction, the classical action of an electron in an external electromagnetic field described by the magnetic vector potential [^5] Note that $\vec{B}=\nabla\times\vec{A}$. Since we only consider the case that $\vec{B}$ is along the $z$-axis, $\nabla\times\vec{A}$ only has the $z$ component. $\vec{A}$ and electric scalar potential $\phi$ is given by (see Appendix C of Yang2023 )
@@
Now we need to add the additional term of potential energy $\bar{U}_{+}$ due to spin magnetic interaction into (44)
@@
The derivation of dynamics equations for a spin-down electron is exactly the same, except there is a sign difference for the potential energy term $\bar{U}_{-}$. The resulting Schrödinger equation is
@@
Generalizing the Schrödinger-Pauli equation for an electron in an external magnetic field $\vec{B}$ along arbitrary orientation is possible. We will leave it for future research and instead focus next on the more interesting quantum phenomenon of entanglement.
Figure 4: Randomness of orientation of intrinsic angular momentum from time perspective. If the measurement occurs at a time $t_{m}$ that when the initial orientation is $\theta\in\Theta^{-}$, the measurement outcome is spin-down. On the other hand, if the measurement occurs at a time $t^{\prime}_{m}$ that when the initial orientation $\theta\in\Theta^{+}$, the measurement outcome is spin-up
@@ V.1 Spin Correlation
In order to investigate the entanglement phenomenon between two electron spins, we need to take a closer look at the fluctuations of spin orientations from the time perspective. As shown earlier, when measuring the spin of an electron using an inhomogeneous external magnetic field along any axis, one will obtain two discrete results, spin-up or spin-down along the axis. Assign this axis as $z$-axis. Whether the result is spin-up or spin-down is random, depending on the initial orientation relative to the $z$-axis, described by the angle $\theta$. To calculate the probability of the initial orientation for $\theta\in\Theta^{+}$ or $\theta\in\Theta^{-}$, we need to analyze how the orientation of intrinsic angular momentum fluctuates before an external magnetic field $B_{z}$ is applied. Recall that we denote $\theta\in\Theta^{+}$ as the orientation of angular momentum trending upward with angle $\theta\in[0,\pi/2]$, and $\theta\in\Theta^{-}$ as the orientation trending downward. It is intuitive to assume that it takes a non-zero amount of time for the orientation to change from $\theta\in\Theta^{+}$ to $\theta\in\Theta^{-}$. That is, it takes a non-zero amount of time for the orientation of intrinsic angular momentum to change from trending upward to trending downward, and vice versa. Suppose at $t_{0}$, $\theta\in\Theta^{+}$. At time $t_{1}^{+}$, $\theta$ changes to trending downward, $\theta\in\Theta^{-}$. So the angular momentum stays trending upward for a period of time $\Delta t^{+}_{1}=t_{1}^{+}-t_{0}$. Then the orientation stays trending downward for a duration $\Delta t^{-}_{1}$ till at time $t_{1}^{-}$, it moves back to point upward. After the orientation stays trending upward for a period of time $\Delta t^{+}_{2}$, it changes to point downward for a period of time $\Delta t^{-}_{2}$. The process of switching the orientation directions continues till $t=\Delta T$, as shown in Figure 4. Statistically, $\{\Delta t^{+}_{i},i\in\mathbb{N}\}$ forms a random variable, and similarly for $\{\Delta t^{+}_{-},i\in\mathbb{N}\}$. Define
@@
we have $\Delta T^{+}+\Delta T^{-}=\Delta T$. Suppose a spin measurement experiment is conducted at $t_{m}$. If $t_{m}$ occurs at the moment when $\theta\in\Theta^{+}$, that is, $t_{m}$ falls into one of the time periods $\Delta t^{+}_{i}$ as depicted in Figure 4, the measurement outcome will be spin-up. On the other hand, if $t_{m}$ falls into one of the time periods $\Delta t^{-}_{i}$, the measurement outcome will be spin-down. Since $\Delta t^{+}_{i}$ and $\Delta t^{-}_{i}$ are random variables, the measurement outcome are random as well, and the probabilities of orientation trending upward and downward are
@@
respectively [^6] Eq. (57) can be considered as the frequency interpretation of $\varrho(\cdot)$, while Eqs. (18) and (19) are the classical interpretation. Both interpretations are considered equivalent.. For an intrinsic angular momentum with completely random orientations, we must have $\Delta T^{+}=\Delta T^{-}$, and therefore, $\varrho(\Theta^{+})=\varrho(\Theta^{-})=1/2$.
The random variable $\{\Delta t^{+}_{i}\geq 0,i\in\mathbb{N}\}$ can follow a probability distribution. Here we will not investigate the actual probability distribution. What is relevant to our investigation is the mean value, denoted as
@@
In summary, the spin model for an electron presented here has the following characteristics:
@@
  • For a particular measurement instant occurred at $t_{m}$, whether the outcome is spin-up or spin-down depends on $\theta\in\Theta^{+}$ or $\theta\in\Theta^{-}$ at $t_{m}$, respectively.
Figure 5: Spin entanglement of two electrons from time perspective. Here, the orientation of intrinsic angular momentum for each electron is completely random, but orientations for both $A$ and $B$ are always opposite to each other. If Alice performs the measurement at a time $t_{m}^{A}$ that when the initial orientation is $\theta_{A}\in\Theta^{-}$, the measurement outcome is spin-down. If Bob performs his measurement at $t_{m}^{B}<t_{1}^{-}$, he will obtain spin-up, as predicted by the singlet spin state. On the other hand, If Bob performs his measurement at $t_{m}^{B^{\prime}}>t_{1}^{-}$ the measurement outcome can be spin-down or spin-up.
@@
Next we want to see what happens when a measurement event occurs. There are two cases here, measurements on $A$ and $B$ occur at the same time, or at different times. First, suppose the two electrons are measured at the same time, which is the usual experiment setup for a typical Bell test. For electron $A$, the probability of measurement outcome can be described as
@@
For electron $B$, because of the correlation $\theta_{B}+\theta_{A}=\pi$, the measurement results will be exactly anti-correlated to the results of $A$. For example, if the measurement result of $A$ is spin-up, it implies the initial condition $\theta_{A}\in\Theta^{+}_{A}$, meaning the orientation of angular momentum is trending upward with any angle $\theta_{A}\in[0,\pi/2)$ relative to the $z$-axis. Since $\theta_{B}+\theta_{A}=\pi$, $\theta_{B}$ must be also trending downward relative to the same direction. Thus, the measurement outcome must be spin-down. Similarly, if the measurement result of $A$ is spin-down, then the measurement result of $B$ must be spin-up. But note that the measurement outcome of $A$ is random, depending on whether the initial condition $\theta_{A}$ is trending upward or downward relative to the $z$-axis at $t_{m}$. The correlation of measurement results can be described by a joint probability density
@@
Before considering the more complicated case that measurements of $A$ and $B$ occur at different times, we need to examine how the correlation $\theta_{B}+\theta_{A}=\pi$ is maintained when the two electrons are separated. Recall that in our spin model, the electron intrinsic angular momentum is not due to rotation as a rigid body, but due to the rotational components of vacuum fluctuations of the electrons, as shown in Appendix B. When the two electrons move away from each other, their motions can be translational such that they have no impact on the orientations of intrinsic angular momenta. Or their motions can have the same rotational effects as well such that the impacts on the orientations of the intrinsic angular momenta are the same. Either case, the correlation $\theta_{B}+\theta_{A}=\pi$ is preserved when the two electrons are physically moved apart. In the modern Bell test experiment Hensen , the two electrons are already separated remotely before being prepared to be entangled. The electron entanglement is achieved by entanglement swapping with photons Hensen . Thus, there is no need to worry about the impacts on the correlation $\theta_{B}+\theta_{A}=\pi$ due to the movement of electrons. Nevertheless, even if $\theta_{B}+\theta_{A}=\pi$ is preserved before measurement, if the measurements on $A$ and $B$ are performed at different times, there is a possibility that the entanglement effect can be diminished. Let’s consider this more subtle situation next.
@@
Suppose the electron pair, with the correlation $\theta_{B}+\theta_{A}=\pi$ established, are separated far away. Electron $A$ is with observer Alice, while electron $B$ is with observer Bob. The correlation $\theta_{B}+\theta_{A}=\pi$ is maintained even though the electron pair are space-like separated. At time $t_{A}^{m}$, Alice performs measurement on $A$. The probability density is given by (59). Alice will get the result randomly. Suppose that she obtains spin-down. At this point, from Bob’s point of view, if he measures electron $B$ along any direction, he will obtain spin-up or spin-down randomly. Now Alice sends Bob her measurement results. The information on the measurement direction is critical. With the information from Alice, Bob infers that $\theta_{A}$ is trending downward relative the measurement direction, and therefore $\theta_{B}$ is trending upward relative the measurement direction due to the correlation $\theta_{B}+\theta_{A}=\pi$. Consequently, he predicts he will obtain spin-up if he perform the measurement on $B$ along the same orientation as Alice. However, due to fluctuation, $\theta_{B}$ can move to trending downward after some time. Bob’s measurement must be performed at a time $t_{B}^{m}$ before $\theta_{B}$ becomes tending downward. This constraint $t_{B}^{m}<t_{1}^{-}$ can be approximately expressed as
@@
Eq. (61) implies that Bob’s prediction on his measurement outcome of spin-up for $B$ is valid only for a period of time. Even though the entanglement between the electron pair is preserved after they are remotely separated, once Alice performs her measurement on $A$, she must send the result to Bob and Bob must confirm his prediction within the time constraint (61). This is different from the prediction of standard quantum mechanics where there is no such time constraint. If $\tau_{B}^{+}$ is sufficiently large, our model will practically give the same prediction as standard quantum mechanics.
One may argue that if $t_{B}^{m}>t_{A}^{m}+\tau_{B}^{+}$, Bob will obtain measurement results of spin-down or spin-up, what is the difference between an entangled pair and a non-entangled electron pair? The answer is that for non-entangled electron pairs, there is no correlation between the measurement results of Alice and Bob. Bob cannot make any prediction on his measurement outcome of $B$ even after he receives the information from Alice.
The analysis here shows that Alice measurements on $A$ will obtain outcome of either spin-up or spin-down randomly, but once the measurement outcome is known, she can predict Bob’s measurement outcome on the remote electron $B$ along the same measurement direction, even though her measurement outcome does not cause physical impact on $B$. There is no mysterious “spooky action in the remote”. The root of the entanglement phenomenon is that the random variables $\theta_{A}$ and $\theta_{B}$ are correlated. Furthermore, these random variables are intrinsically local to the electron, it is possible to maintain the correlation even though the two electrons are remotely separated.
V.2 Singlet States
If the correlation for the entangled electron pair is described by $\theta_{A}+\theta_{B}=\pi$ instead, the measurement outcome on $A$ will be exactly opposite to the measurement outcome on $B$. The joint probability density is given by
In the case that the initial orientations of angular momentum of electron $A$ are uniformly distributed for all $\theta\in[0,\pi]$, one has $\varrho(\Theta^{+}_{A})=\varrho(\Theta^{-}_{A})=1/2$, (62) becomes
where the superscript $z$ is added to explicitly show the context dependency on the $z$-axis. However, (63) itself does not uniquely specify to the effect of correlation when the angular momentum orientations of the electron pair are always opposite to each other, because (63) can correspond to two possible correlations, as shown in Figure 6. To uniquely specify the effect of correlation when the angular momentum orientations of the electron pair are always opposite to each other, one needs to show the probability density along another axis perpendicular $z$-axis, say, the $y$-axis. Consider the case that the correlations between the electron pair are such that not only $\theta_{A}^{z}+\theta^{z}_{B}=\pi$ along the $z$-axis, but also $\theta_{A}^{y}+\theta^{y}_{B}=\pi$ along the $y$-axis. This implies that we must also have
Figure 6: Correlation of intrinsic angular momenta between two electrons. a.) Correlation corresponds to the singlet state $|\Psi^{-}\rangle$; b.) Correlation corresponds to Bell state $|\Psi^{+}\rangle$. Note that for both a.) and b.), we have $\theta_{A}^{z}+\theta^{z}_{B}=\pi$. It shows $\theta_{A}^{z}+\theta^{z}_{B}=\pi$ alone is unable to distinguish the correlations between a.) and b.).
When a pair of electrons are described by both (63) and (64), the correlation between the two electrons is equivalent to that is described by the spin singlet state in standard quantum mechanics.
We will further verify the equivalence in the next section when calculating the Bell-CHSH inequality.

### v3

@@ III.3 Subsequent Measurement with a Rotated Stern-Gerlach Apparatus
In Appendix D, it is verified that (36) is exactly the same result predicted by standard quantum mechanics.
@@ IV Derivation of the Schrödinger-Pauli Equation
The framework to derive the law of dynamics for an electron in a external magnetic field $B_{z}$, for a cumulative period from $t_{a}$ to $t_{b}$, is similar to that described in Appendix A, except that we need to add a new term in the Lagrangian due to the interaction of spin and the external magnetic field. Based on the results shown in (16) and (29), the probability density can be written as
@@
$\begin{split}\overline{U}&=-\int p(\theta)\vec{\mu}\cdot\vec{B}d\theta\\ &=\frac{e\hbar}{2m}\int p(\theta)B_{z}\cos(\theta)d\theta,\end{split}$ (38)
@@
$\begin{split}\overline{U}&=\frac{e\hbar}{2m}\int\rho(\mathbf{x},t,\theta)B_{z}\cos(\theta)d\theta d^{3}\mathbf{x}\\ &=\frac{e\hbar}{2m}\sum_{i}\int\sigma_{i}(\mathbf{x},t,\theta_{i})\delta(\theta-\theta_{i})B_{z}\cos(\theta)d\theta d^{3}\mathbf{x}.\end{split}$ (40)
@@
$\begin{split}\overline{U}&=\frac{e\hbar}{2m}\{\int\sigma_{0}(\mathbf{x},t,0)\delta(\theta)\ B_{z}\cos(\theta)d\theta d^{3}\mathbf{x}\\ &+\int\sigma_{1}(\mathbf{x},t,\pi)\delta(\theta-\pi)B_{z}\cos(\theta)d\theta d^{3}\mathbf{x}\}\\ &=\frac{e\hbar}{2m}\{\int\sigma_{0}(\mathbf{x},t,0)B_{z}d^{3}\mathbf{x}-\int\sigma_{1}(\mathbf{x},t,\pi)B_{z}d^{3}\mathbf{x}\}.\end{split}$ (41)
Here we run into a problem. In the Stern-Gerlach experiment, it is confirmed that electrons with spin-up follow a different trajectory from that of electrons with spin-down. An electron with either spin-up or spin-down follows separated laws of dynamics. We cannot simply add the two terms in (41) together. Instead, a more proper notation should treat the probability density as a two-component vector. Thus,
$\overline{U}=\frac{e\hbar}{2m}\int B_{z}\left(\begin{array}[]{c}\sigma_{0}(\mathbf{x},t,0)\\ -\sigma_{1}(\mathbf{x},t,\pi)\end{array}\right)dx:=\left(\begin{array}[]{c}\overline{U}_{+}\\ \overline{U}_{-}\end{array}\right).$ (42)
For further simplification of notation, denote $\rho_{+}(x,t)=\sigma_{0}(x,t,0)$ and $\rho_{-}(x,t)=\sigma_{1}(x,t,\pi)$. They should satisfy the normalization condition
$\int[\rho_{+}(\mathbf{x},t)+\rho_{-}(\mathbf{x},t)]dx=1.$ (43)
First we will derive the law of dynamics for the electron with spin up. Without considering the contribution from the spin magnetic interaction, the classical action of an electron in an external electromagnetic field described by the magnetic vector potential [^5] Note that $\vec{B}=\nabla\times\vec{A}$. Since we only consider the case where $\vec{B}$ is along the $z$-axis, $\nabla\times\vec{A}$ only has the $z$ component. $\vec{A}$ and electric scalar potential $\phi$ is given by (see Appendix C of Yang2023 )
@@
Now we need to add the additional term of potential energy $\overline{U}_{+}$ due to spin magnetic interaction into (44)
@@
The derivation of the dynamics equations for a spin-down electron is exactly the same, except there is a sign difference for the potential energy term $\overline{U}_{-}$. The resulting Schrödinger equation is
@@
Generalizing the Schrödinger-Pauli equation for an electron in an external magnetic field $\vec{B}$ along arbitrary orientation is the natural next step, as it will demonstrate the rotational property of spin. We will leave it for future research.
Figure 4: Randomness of orientation of intrinsic angular momentum from time perspective. If the measurement occurs at a time $t_{m}$ that the initial orientation is $\theta\in\Theta^{-}$, the measurement outcome is spin-down. On the other hand, if the measurement occurs at a time $t^{\prime}_{m}$ that the initial orientation $\theta\in\Theta^{+}$, the measurement outcome is spin-up
@@ V.1 Spin Correlation
To investigate the entanglement phenomenon between two electron spins, we need to take a closer look at the time dynamics of the fluctuations of spin orientations. As shown earlier, when measuring the spin of an electron using an inhomogeneous external magnetic field along any axis, one will obtain two discrete results, spin-up or spin-down along the axis. Assign this axis as the $z$-axis. Whether the result is spin-up or spin-down is random, depending on the initial orientation relative to the $z$-axis, described by the angle $\theta$. To calculate the probability of the initial orientation for $\theta\in\Theta^{+}$ or $\theta\in\Theta^{-}$, we need to analyze how the orientation of intrinsic angular momentum fluctuates before an external magnetic field $B_{z}$ is applied. Recall that we denote $\theta\in\Theta^{+}$ as the orientation of the angular momentum that is trending upward with angle $\theta\in[0,\pi/2]$, and $\theta\in\Theta^{-}$ as the orientation that is trending downward. The amount of time for the orientation to change from $\theta\in\Theta^{+}$ to $\theta\in\Theta^{-}$ must be non-zero. That is, it takes a non-zero amount of time for the orientation of intrinsic angular momentum to change from trending upward to trending downward, and vice versa. Suppose that at $t_{0}$, $\theta\in\Theta^{+}$. At time $t_{1}^{+}$, $\theta$ changes to a downward trend, $\theta\in\Theta^{-}$. So, the angular momentum continues to trend upward for a period of time $\Delta t^{+}_{1}=t_{1}^{+}-t_{0}$. Then the orientation stays trending downward for a duration $\Delta t^{-}_{1}$ until at time $t_{1}^{-}$, it moves back to pointing upward. After the orientation continues to trend upward for a period of time $\Delta t^{+}_{2}$, it changes to point downward for a period of time $\Delta t^{-}_{2}$. The process of switching orientation directions continues to $t=\Delta T$, as shown in Figure 4. Statistically, $\{\Delta t^{+}_{i},i\in\mathbb{N}\}$ forms a random variable, and so as for $\{\Delta t^{-}_{i},i\in\mathbb{N}\}$. Define
@@
we have $\Delta T^{+}+\Delta T^{-}=\Delta T$. Suppose that a spin measurement experiment is conducted at $t_{m}$. If $t_{m}$ occurs at the moment when $\theta\in\Theta^{+}$, that is, $t_{m}$ falls into one of the time periods $\Delta t^{+}_{i}$, as shown in Figure 4, the measurement result will be spin-up. On the other hand, if $t_{m}$ falls into one of the time periods $\Delta t^{-}_{i}$, the measurement result will be spin-down. Since $\Delta t^{+}_{i}$ and $\Delta t^{-}_{i}$ are random variables, the measurement outcome are random as well, and the probabilities of orientation trending upward and downward are
@@
respectively [^6] Eq. (57) can be considered as the frequency interpretation of $\varrho(\cdot)$ in Eqs. (18) and (19), while Eqs. (18) and (19) are the classical interpretation. Both interpretations are considered equivalent.. For an intrinsic angular momentum with completely random orientations, we must have $\Delta T^{+}=\Delta T^{-}$, and therefore $\varrho(\Theta^{+})=\varrho(\Theta^{-})=1/2$.
The random variable $\{\Delta t^{+}_{i}\geq 0,i\in\mathbb{N}\}$ can follow a probability distribution. Here we will not investigate the actual probability distribution. What is relevant to our investigation is the expectation value, denoted as
@@
In summary, the spin model of an electron developed here has the following characteristics:
@@
  • For a particular measurement event occurred at $t_{m}$, whether the outcome is spin-up or spin-down depends on $\theta\in\Theta^{+}$ or $\theta\in\Theta^{-}$ at $t_{m}$, respectively.
@@
Next, we want to see what happens when a measurement event occurs. There are two cases here; measurements on $A$ and $B$ occur at the same time or at different times. First, suppose that the two electrons are measured at the same time, which is the usual experiment setup for a typical Bell test. For electron $A$, the probability of measurement outcome can be described as
@@
For electron $B$, due to the correlation $\theta_{B}+\theta_{A}=\pi$, the measurement results will be exactly anti-correlated to the results of $A$. For example, if the measurement result of $A$ is spin-up, it implies the initial condition $\theta_{A}\in\Theta^{+}_{A}$, which means that the orientation of the angular momentum is trending upward with any angle $\theta_{A}\in[0,\pi/2)$ relative to the $z$-axis. Since $\theta_{B}+\theta_{A}=\pi$, $\theta_{B}$ must also trend downward relative to the same direction. Thus, the measurement outcome must be spin-down. Similarly, if the measurement result of $A$ is spin-down, then the measurement result of $B$ must be spin-up. However, note that the measurement result of $A$ is random, depending on whether the initial condition $\theta_{A}$ is trending upward or downward relative to the $z$-axis at $t_{m}$. The correlation of measurement results can be described by a joint probability density
@@
where the superscript $z$ is added to explicitly show the context dependency on the $z$-axis. However, Eq. (61) itself does not uniquely specify the effect of correlation when the angular momentum orientations of the electron pair are always opposite to each other, because (61) can correspond to two possible correlations, as shown in Figure 5. To uniquely specify the effect of correlation when the angular momentum orientations of the electron pair are always opposite to each other, one needs to show the probability density along another axis perpendicular $z$-axis, say the $y$-axis. Consider the case that the correlations between the electron pair are such that not only $\theta_{A}^{z}+\theta^{z}_{B}=\pi$ along the $z$-axis, but also $\theta_{A}^{y}+\theta^{y}_{B}=\pi$ along the $y$-axis. This implies that we must also have
Figure 5: Correlation of intrinsic angular momenta between two electrons. a.) Correlation corresponds to the singlet state $|\Psi^{-}\rangle$; b.) Correlation corresponds to Bell state $|\Psi^{+}\rangle$. Note that for both a.) and b.), we have $\theta_{A}^{z}+\theta^{z}_{B}=\pi$. It shows $\theta_{A}^{z}+\theta^{z}_{B}=\pi$ alone is unable to distinguish the correlations between a.) and b.). Thus, we need to also specify the correlation along the $y$-axis.
When a pair of electrons is described by both (61) and (62), the correlation between the two electrons is equivalent to that described by the spin singlet state in standard quantum mechanics.
The equivalence of (63) versus (61) and (62) will be verified in the next section when we calculate the Bell-CHSH inequality.
@@
Before considering the more complicated case where measurements of $A$ and $B$ occur at different times, we need to clarify two important questions. First, how is the correlation $\theta_{B}+\theta_{A}=\pi$ maintained when the two electrons are separated? Recall that in our spin model, the intrinsic angular momentum of the electron is not due to rotation as a rigid body, but due to the rotational components of vacuum fluctuations of the electrons, as shown in Appendix B. When the two electrons move away from each other, their motions can be translational such that they have no impact on the orientations of intrinsic angular momenta. Or, their motions can have the same rotational effects on both electrons such that the impacts on the orientations of the intrinsic angular momenta are the same. In either case, the correlation $\theta_{B}+\theta_{A}=\pi$ is preserved when the two electrons physically move apart. In the modern Bell test experiment Hensen , the two electrons have already been remotely separated before being prepared to be entangled. The electron entanglement is achieved by entanglement swapping with photons Hensen . Thus, there is no need to worry about the impacts on the correlation $\theta_{B}+\theta_{A}=\pi$ due to the movement of electrons.
Second, how is the spin correlation different from a classical correlation? An example of a classical correlation is a pair of shoes placed in two sealed boxes. The two boxes are remotely separated. When an observer opens one box and finds that the shoe is for the left foot, she will immediately predict that the other shoe located remotely is for the right foot. The spin correlation presented here has two important differences from this classical correlation. Firstly, when an observer measures the spin of electron A, the result is completely random. This randomness is intrinsic, while the left or right property of the shoe in the classical correlation example is a deterministic property. The properties correlated in a pair of entangled spins are the spin orientations, and they are random variables. Secondly, to uniquely specify the correlation between two random variables of orientation, one needs to specify the correlation on both the $z$-axis and the $y$-axis, as shown in (61) and (62).
V.2 Time Dependency of Measurement Results on A Pair of Entangled Spins
The spin entanglement described in our model implies an interesting property. In the singlet spin state, even if $\theta_{B}+\theta_{A}=\pi$ is preserved before measurement, if the measurements on $A$ and $B$ are performed at different times, there is a possibility that the entanglement effect can be diminished. Let us next examine this more subtle situation.
@@ V.2 Time Dependency of Measurement Results on A Pair of Entangled Spins
Suppose that the electron pairs with the correlation $\theta_{B}+\theta_{A}=\pi$ established, are separated far away. Electron $A$ is with observer Alice, while electron $B$ is with observer Bob. The correlation $\theta_{B}+\theta_{A}=\pi$ is maintained even though the electron pair are space-like separated. At time $t_{A}^{m}$, Alice performs measurement on $A$. The probability density is given by (59). Alice will obtain the measurement result randomly. Suppose that she observes spin-down. At this point, from Bob’s point of view, if he measures electron $B$ along any direction, he will obtain spin-up or spin-down randomly. Now Alice sends Bob her measurement results along with information on the direction of measurement. With Alice’s information, Bob infers that $\theta_{A}$ trendes downward relative to the measurement direction, and therefore $\theta_{B}$ trendes upward relative to the same measurement direction due to the correlation $\theta_{B}+\theta_{A}=\pi$. Consequently, he predicts that he will obtain spin-up if he performs the measurement on $B$ along the same orientation as Alice. However, due to fluctuation, $\theta_{B}$ can become trending downward after some time. Bob’s measurement must be performed at a time $t_{B}^{m}$ before $\theta_{B}$ becomes tending downward. This constraint can be approximately expressed as
@@
Eq. (64) implies that Bob’s prediction on his measurement outcome of spin-up for $B$ is valid only for a period of time. Although the entanglement between the electron pairs is preserved after they are remotely separated, once Alice takes her measurement in $A$, Bob must perform his measurement on $B$ within the time constraint (64). This is different from the prediction of standard quantum mechanics, where there is no such time constraint. If $\tau_{B}^{+}$ is sufficiently large, our model will practically give the same prediction as standard quantum mechanics.
One may argue that if $t_{B}^{m}>t_{A}^{m}+\tau_{B}^{+}$, Bob will obtain measurement results of spin-down or spin-up, what is the difference between an entangled pair and a non-entangled electron pair? The answer is that for non-entangled electron pairs, there is no correlation between the measurement results of Alice and Bob. Bob cannot make any prediction on the outcome of his measurement of $B$ even immediately after receiving the information from Alice.
This interesting property of time dependency of the measurement results on a pair of entangled spin will be used to design a modified Bell test in Section VI.2.
Figure 6: Spin entanglement of two electrons from time perspective. Here, the orientation of intrinsic angular momentum for each electron is completely random, but the orientations for both $A$ and $B$ are always opposite to each other. If Alice performs the measurement at time $t_{m}^{A}$ that the initial orientation is $\theta_{A}\in\Theta^{-}$, the measurement outcome is spin-down. If Bob performs his measurement at $t_{m}^{B}<t_{1}^{-}$, he will obtain spin-up, as predicted by the singlet spin state. On the other hand, if Bob performs his measurement at $t_{m}^{B^{\prime}}>t_{1}^{-}$ the measurement outcome can be spin-down or spin-up.